Hindustan Book Agency

Conference on Number Theory and Discrete Mathematics in honour of Srinivasa Ramanujan in Chandigarh from October 2-6,2000.

Multiple Polylogarithms: An Introduction

by

M. WALDSCHMIDT (^{∗})

Dedicated to Professor R.P. Bambah on his 75th birthday

Abstract

Multiple polylogarithms in a single variable are defined by

Li_{(s_{1},...,s_{k})}(z) = ∑_{n_{1}>n_{2}>···>n_{k}≥1} \frac{z^{n_{1}}}{n_{1}^{s_{1}} ⋯ n_{k}^{s_{k}}} ,

when s_{1}, …, s_{k} are positive integers and z a complex number in the unit disk. For k = 1, this is the classical polylogarithm Li_{s}(z). These multiple polylogarithms can be defined also in terms of iterated Chen integrals and satisfy shuffle relations. Multiple polylogarithms in several variables are defined for s_{i} ≥ 1 and |z_{i}| < 1

(1 ≤ i ≤ k) by

Li_{(s_{1},...,s_{k})}(z_{1}, …, z_{k}) = ∑_{n_{1}>n_{2}>···>n_{k}≥1} \frac{z_{1}^{n_{1}} ⋯ z_{k}^{n_{k}}}{n_{1}^{s_{1}} ⋯ n_{k}^{s_{k}}} ,

and they satisfy not only shuffle relations, but also stuffle relations. When one specializes the shuffle relations in one variable at z = 1 and the stuffle relations in several variables at z_{1} = ⋯ = z_{k} = 1, one gets linear or quadratic dependence relations between the Multiple Zeta Values

ζ(s_{1}, …, s_{k}) = ∑_{n_{1}>n_{2}>···>n_{k}≥1} \frac{1}{n_{1}^{s_{1}} ⋯ n_{k}^{s_{k}}}

which are defined for k, s_{1}, …, s_{k} positive integers with s_{1} ≥ 2. The Main Diophantine Conjecture states that one obtains in this way all algebraic relations between these MZV.

Classification AMS: 11J91, 33E30

(^{∗}) Institut de Mathématiques, Université P. et M. Curie (Paris VI), Théorie des Nombres Case 247, F–75013 PARIS miw@math.jussieu.fr

    http://www.math.jussieu.fr/∼miw/

0. Introduction

A long term project is to determine all algebraic relations among the values

π, ζ(3), ζ(5), …, ζ(2n + 1), …

of the Riemann zeta function

ζ(s) = ∑_{n≥1} \frac{1}{n^{s}} ·

So far, one only knows that the first number in this list, π, is transcendental, that the second one, ζ(3), is irrational, and that the other ones span a ℚvector space of infinite dimension [Ri1], [BR]. See also [Ri2], [Zu1] and [Zu2].

The expected answer is disappointingly simple: it is widely believed that there are no relations, which means that these numbers should be algebraically independent: (?) For any n ≥ 0 and any nonzero polynomial P ∈ ℤ[X_{0}, …, X_{n}],

P (π, ζ(3), ζ(5), …, ζ(2n + 1)) ≠ 0.

If true, this property would mean that there is no interesting algebraic structure.

The situation changes drastically if we enlarge our set so as to include the so-called Multiple Zeta Values (MZV, also called Euler-Zagier numbers or Polyzêta – see [Eu], [Z] and [C]):

ζ(s_{1}, …, s_{k}) = ∑_{n_{1}>n_{2}>···>n_{k}≥1} \frac{1}{n_{1}^{s_{1}} ⋯ n_{k}^{s_{k}}},

which are defined for k, s_{1}, …, s_{k} positive integers with s_{1} ≥ 2. It may be hoped that the initial goal would be reached if one could determine all algebraic relations between the MZV. Now there are plenty of relations between them, providing a rich algebraic structure. One type of such relations arises when one multiplies two such series: it is easy to see that one gets a linear combination of MZV. There is another type of algebraic relations between MZV, coming from their expressions as integrals. Again the product of two such integrals is a linear combination of MZV. Following [B^{3}], we will use the name stuffle for the relations arising from the series, and shuffle for those arising from the integrals.

The Main Diophantine Conjecture (Conjecture 5.3 below) states that these relations are sufficient to describe all algebraic relations between MZV. One should be careful when stating such a conjecture: it is necessary to include some relations which are deduced from the stuffle and shuffle relations applied to divergent series (i.e. with s_{1} = 1).

There are several ways of dealing with the divergent case. Here, we use the multiple polylogarithms

Li_{(s_{1},...,s_{k})}(z) = ∑_{n_{1}>n_{2}>···>n_{k}≥1} \frac{z^{n_{1}}}{n_{1}^{s_{1}} ⋯ n_{k}^{s_{k}}},

which are defined for |z| < 1 when s_{1}, …, s_{k} are all ≥ 1, and which are also defined for |z| = 1 if s_{1} ≥ 2.

These multiple polylogarithms can be expressed as iterated Chen integrals, and from this representation one deduces shuffle relations. There is no stuffle relations for multiple polylogarithms in a single variable, but one recovers them by introducing the multivariables functions

Li_{(s_{1},...,s_{k})}(z_{1}, …, z_{k}) = ∑_{n_{1}>n_{2}>···>n_{k}≥1} \frac{z_{1}^{n_{1}} ⋯ z_{k}^{n_{k}}}{n_{1}^{s_{1}} ⋯ n_{k}^{s_{k}}}  (s_{i} ≥ 1, 1 ≤ i ≤ k)

which are defined not only for |z_{1}| < 1 and |z_{i}| ≤ 1 (2 ≤ i ≤ k), but also for |z_{i}| ≤ 1 (1 ≤ i ≤ k) if s_{1} ≥ 2.

1. Multiple Polylogarithms in One Variable and Multiple Zeta Values

Let k, s_{1}, …, s_{k} be positive integers. Write \underline{s} in place of (s_{1}, …, s_{k}). One defines a complex function of one variable by

Li\underline{s}(z) = ∑_{n_{1}>n_{2}>···>n_{k}≥1} \frac{z^{n_{1}}}{n_{1}^{s_{1}} ⋯ n_{k}^{s_{k}}} ·

This function is analytic in the open unit disk, and, in the case s_{1} ≥ 2, it is also continuous on the closed unit disk. In the latter case we have

ζ(\underline{s}) = Li\underline{s}(1).

One can also define in an equivalent way these functions by induction on the number p = s_{1} + ⋯ + s_{k} (the weight of \underline{s}) as follows. Plainly we have

(1.1)  z \frac{d}{dz} Li_{(s_{1},...,s_{k})}(z) = Li_{(s_{1}−1,s_{2},...,s_{k})}(z) if s_{1} ≥ 2

and

(1.2)  (1 − z)\frac{d}{dz} Li_{(1,s_{2},...,s_{k})}(z) = Li_{(s_{2},...,s_{k})}(z) if k ≥ 2.

Together with the initial conditions

(1.3)  Li\underline{s}(0) = 0,

the differential equations (1.1) and (1.2) determine all the Li\underline{s}.

Therefore, as observed by M. Kontsevich (cf. [Z]; see also [K] Chap. XIX, § 11 for an early reference to H. Poincaré, 1884), an equivalent definition for Li\underline{s} is given by integral formulae as follows. Starting(^{∗}) with k = s = 1, we write

Li_{1}(z) = − log(1 − z) = ∫_{0}^{z} \frac{dt}{1 − t},

where the complex integral is over any path from 0 to z inside the unit circle. From the differential equations (1.1) one deduces, by induction, for s ≥ 2,

Li_{s}(z) = ∫_{0}^{z} Li_{s−1}(t)\frac{dt}{t} = ∫_{0}^{z} \frac{dt_{1}}{t_{1}} ∫_{0}^{t_{1}} \frac{dt_{2}}{t_{2}} ⋯ ∫_{0}^{t_{s−2}} \frac{dt_{s−1}}{t_{s−1}} ∫_{0}^{t_{s−1}} \frac{dt_{s}}{1 − t_{s}} ·

In the last formula, the complex integral over t_{1}, which is written on the left (and which is the last one to be computed), is over any path inside the unit circle from 0 to z, the second one over t_{2} is from 0 to t_{1},…, and the last one over t_{s} on the right, which is the first one to be computed, is from 0 to t_{s−1}.

Chen iterated integrals (see [K] Chap. XIX, § 11) provide a compact form for such expressions as follows. For φ_{1}, …, φ_{p} differential forms and x, y complex numbers, define inductively

∫_{x}^{y} φ_{1} ⋯ φ_{p} = ∫_{x}^{y} φ_{1}(t) ∫_{x}^{t} φ_{2} ⋯ φ_{p}.

For \underline{s} = (s_{1}, …, s_{k}), set

ω\underline{s} = ω_{0}^{s_{1}−1}ω_{1} ⋯ ω_{0}^{s_{k}−1}ω_{1},

where

ω_{0}(t) = \frac{dt}{t} and ω_{1}(t) = \frac{dt}{1 − t}·

(^{∗}) This induction could as well be started from k = 0, provided that we set Li_{∅}(z) = 1.

Then the differential equations (1.1) and (1.2) with initial conditions (1.3) can be written

(1.4)  Li\underline{s}(z) = ∫_{0}^{z} ω\underline{s}.

Example. Given a string a_{1}, …, a_{k} of integers, the notation {a_{1}, …, a_{k}}_{n} stands for the kn-tuple

(a_{1}, …, a_{k}, …, a_{1}, …, a_{k}),

where the string a_{1}, …, a_{k} is repeated n times.

For any n ≥ 1 and |z| < 1 we have

(1.5)  Li_{\{1\}_{n}}(z) = \frac{1}{n!}(log(1/(1 − z)))^{n},

which can be written in terms of generating series as

∑_{n=0}^{∞} Li_{{1}_{n}}(z)x^{n} = (1 − z)^{−x}.

The constant term Li_{{1}_{0}}(z) is 1.

2. Shuffle Product and the First Standard Relations

Denote by X = {x_{0}, x_{1}} the alphabet with two letters and by X^{∗} the set of words on X. A word is nothing else than a non-commutative monomial in the two letters x_{0} and x_{1}. The linear combinations of such words with rational coefficients

∑_{u} c_{u}u,

where {c_{u} ; u ∈ X^{∗}} is a set of rational numbers with finite support, is the non-commutative ring ℌ = ℚ⟨x_{0}, x_{1}⟩. The product is concatenation, the unit is the empty word e. We are interested with the set X^{∗}x_{1} of words which end with x_{1}. The linear combinations of such words is a left ideal of ℌ which we denote by ℌx_{1}. Also we denote by ℌ^{1} the subalgebra ℚe+ℌx_{1} of ℌ.

For s a positive integer, set y_{s} = x_{0}^{s−1}x_{1}. Next, for a tuple \underline{s} = (s_{1}, …, s_{k}) of positive integers, define y\underline{s} = y_{s_{1}} ⋯ y_{s_{k}}. Hence the set X^{∗}x_{1} is also the set of words y\underline{s}, where k, s_{1}, …, s_{k} run over the set of all positive integers. We define L̂i_{u}(z) for u ∈ X^{∗}x_{1} by L̂i_{u}(z) = Li\underline{s}(z) when u = y\underline{s}. By linearity we extend the definition of L̂i_{u}(z) to ℌ^{1}:

L̂i_{v}(z) = ∑_{u} c_{u}L̂i_{u}(z) for v = ∑_{u} c_{u}u

where u ranges over a finite subset of {e}∪X^{∗}x_{1} and c_{u} ∈ ℚ, while L̂i_{e}(z) = 1. The set of convergent words is the set, denoted by {e}∪x_{0}X^{∗}x_{1}, of words which start with x_{0} and end with x_{1} together with the empty word e. The ℚ-vector subspace they span in ℌ is the subalgebra ℌ^{0} = ℚe + x_{0}ℌx_{1} of ℌ^{1}, and for v in ℌ^{0} we set

ζ̂(v) = L̂i_{v}(1)

so that ζ̂: ℌ^{0} → ℝ is a ℚ-linear map and

ζ̂(y\underline{s}) = ζ(\underline{s})

for y\underline{s} in x_{0}X^{∗}x_{1}.

Definition. The shuffle product of two words in X^{∗} is the element in ℌ which is defined inductively as follows:

e� u = u� e = u

for any u in X^{∗}, and

(x_{i}u)� (x_{j}v) = x_{i}(u� (x_{j}v)) + x_{j}((x_{i}u)� v)

for u, v in X^{∗} and i, j equal to 0 or 1.

This product is extended by distributivity with respect to the addition to ℌ and defines a commutative and associative law. Moreover ℌ^{0} and ℌ^{1} are stable under � . We denote by ℌ�^{0} ⊂ ℌ�^{1} ⊂ ℌ_{�} the algebras where the underlying sets are ℌ^{0} ⊂ ℌ^{1} ⊂ ℌ respectively and the product is � . Radford’s Theorem gives the structure of these algebras: they are (commutative) polynomials algebras on the set of Lyndon words (see for instance [R]).

Computing the product L̂i_{u}(z)L̂i_{u′}(z) of the two associated Chen iterated integrals yields (see [MPH], Th. 2):

Proposition 2.1. For u and u′ in ℌ�^{1} ,

L̂i_{u}(z)L̂i_{u′}(z) = L̂i_{u� u′}(z).

For instance from

x_{1}� (x_{0}x_{1}) = x_{1}x_{0}x_{1} + 2x_{0}x_{1}^{2}

we deduce

(2.2)  Li_{1}(z)Li_{2}(z) = Li_{(1,2)}(z) + 2Li_{(2,1)}(z).

Setting z = 1, we deduce from Proposition 2.1:

(2.3)  ζ̂(u)ζ̂(u′) = ζ̂(u� u′)

for u and u′ in ℌ�^{0} .

These are the first standard relations between multiple zeta values.

3. Shuffle Product for Multiple Polylogarithms in Several Variables

The functions of k complex variables(^{∗})

Li\underline{s}(z_{1}, …, z_{k}) = ∑_{n_{1}>n_{2}>···>n_{k}≥1} \frac{z_{1}^{n_{1}} ⋯ z_{k}^{n_{k}}}{n_{1}^{s_{1}} ⋯ n_{k}^{s_{k}}}

have been considered as early as 1904 by N. Nielsen, and rediscovered later by A.B. Goncharov [G1,G2]. Recently, J. Écalle [É] used them for z_{i} roots of unity (in case s_{1} ≥ 2): these are the decorated multiple polylogarithms. Of course one recovers the one variable functions Li\underline{s}(z) by specializing z_{2} = ⋯ = z_{k} = 1. For simplicity we write Li\underline{s}(\underline{z}), where \underline{z} stands for (z_{1}, …, z_{k}). There is an integral formula which extends (1.4). Define

ω_{z}(t) = \{ \begin{matrix}\frac{zdt}{1 − zt} & if z ≠ 0, \\ \frac{dt}{t} & if z = 0. \end{matrix}

(^{∗}) Our notation for

Li_{(s_{1},...,s_{k})}(z_{1}, …, z_{k})

is the same as in [H], [W] or [C], but for Goncharov’s [G2] it corresponds to

Li_{(s_{k},...,s_{1})}(z_{k}, …, z_{1}).

From the differential equations

z_{1} \frac{∂}{∂z_{1}} Li\underline{s}(\underline{z}) = Li_{(s_{1}−1,s_{2},...,s_{k})}(\underline{z}) if s_{1} ≥ 2

and

(1 − z_{1})\frac{∂}{∂z_{1}} Li_{(1,s_{2},...,s_{k})}(\underline{z}) = Li_{(s_{2},...,s_{k})}(z_{1}z_{2}, z_{3}, …, z_{k})

generalizing (1.1) and (1.2), we deduce

Li\underline{s}(\underline{z}) = ∫_{0}^{1} ω_{0}^{s_{1}−1}ω_{z_{1}}ω_{0}^{s_{2}−1}ω_{z_{1}z_{2}} ⋯ ω_{0}^{s_{k}−1}ω_{z_{1}···z_{k}}.

Because of the occurrence of the products z_{1} ⋯ z_{j} (1 ≤ j ≤ k), it is convenient (see for instance [G1] and [B^{3}L]) to perform the change of variables

y_{j} = z_{1}^{−1} ⋯ z_{j}^{−1} (1 ≤ j ≤ k) and z_{j} = \frac{y_{j−1}}{y_{j}} (1 ≤ j ≤ k)

with y_{0} = 1, and to introduce the differential forms

ω′_{y}(t) = −ω_{y^{−1}}(t) = \frac{dt}{t − y},

so that ω′_{0} = ω_{0} and ω′_{1} = −ω_{1}. Also define

λ\binom{s_{1}, …, s_{k}}{y_{1}, …, y_{k}} = Li\underline{s}(1/y_{1}, y_{1}/y_{2}, …, y_{k−1}/y_{k})

= ∑_{ν_{1}≥1} ⋯ ∑_{ν_{k}≥1} ∏_{j=1}^{k} y_{j}^{−ν_{j}} (∑_{i=j}^{k} ν_{i})^{−s_{j}}

= (−1)^{p} ∫_{Δ_{p}} ω_{0}^{s_{1}−1}ω′_{y_{1}} ⋯ ω_{0}^{s_{k}−1}ω′_{y_{k}}.

With this notation some formulae are simpler. For instance the shuffle relation is easier to write with λ: the shuffle is defined on words on the alphabet {ω′_{y} ; y ∈ ℂ}, (including y = 0), inductively by

(ω′_{y}u)� (ω′_{y′}v) = ω′_{y}(u� ω′_{y′}v)) + ω′_{y′}((ω′_{y}u)� v).

4. Stuffle Product and the Second Standard Relations

The functions Li\underline{s}(\underline{z}) satisfy not only shuffle relations, but also stuffle relations arising from the product of two series:

(4.1)  Li\underline{s}(\underline{z})Li\underline{s}′(\underline{z}′) = ∑_{\underline{s}′′} Li\underline{s}′′(\underline{z}′′),

where the notation is as follows: \underline{s}′′ runs over the tuples (s′′_{1}, …, s′′_{k′′}) obtained from \underline{s} = (s_{1}, …, s_{k}) and \underline{s}′ = (s′_{1}, …, s′_{k′}) by inserting, in all possible ways, some 0 in the string (s_{1}, …, s_{k}) as well as in the string (s′_{1}, …, s′_{k′}) (including in front and at the end), so that the new strings have the same length k′′, with max{k, k′} ≤ k′′ ≤ k + k′, and by adding the two sequences term by term. For each such \underline{s}′′, the component z′′_{i} of \underline{z} is z_{j} if the corresponding s′′_{i} is just s_{j} (corresponding to a 0 in \underline{s}′), it is z′_{ℓ} if the corresponding s′′_{i} is s′_{ℓ} (corresponding to a 0 in \underline{s}), and finally it is z_{j}z′_{ℓ} if the corresponding s′′_{i} is s_{j} + s′_{ℓ}. For instance

\begin{matrix}\underline{s} & s_{1} & s_{2} & 0 & s_{3} & s_{4} & ⋯ & 0 \\ \underline{s}′ & 0 & s′_{1} & s′_{2} & 0 & s′_{3} & ⋯ & s′_{k′} \\ \underline{s}′′ & s_{1} & s_{2} + s′_{1} & s′_{2} & s_{3} & s_{4} + s′_{3} & ⋯ & s′_{k′} \\ \underline{z}′′ & z_{1} & z_{2}z′_{1} & z′_{2} & z_{3} & z_{4}z′_{3} & ⋯ & z′_{k′}. \end{matrix}

Of course the 0’s are inserted so that no s′′_{i} is zero.

Examples. For k = k′ = 1 the stuffle relation (4.1) yields

(4.2)  Li_{s}(z)Li_{s′}(z′) = Li_{(s,s′)}(z, z′) + Li_{(s′,s)}(z′, z) + Li_{s+s′}(zz′),

while for k = 1 and k′ = 2 we have

Li_{s}(z)Li_{(s′_{1},s′_{2})}(z′_{1}, z′_{2}) = Li_{(s,s′_{1},s′_{2})}(z, z′_{1}, z′_{2}) + Li_{(s′_{1},s,s′_{2})}(z′_{1}, z, z′_{2})

(4.3) + Li_{(s′_{1},s′_{2},s)}(z′_{1}, z′_{2}, z) + Li_{(s+s′_{1},s′_{2})}(zz′_{1}, z′_{2}) + Li_{(s′_{1},s+s′_{2})}(z′_{1}, zz′_{2}).

The stuffle product ⋆ is defined on X^{∗} inductively by

e ⋆ u = u ⋆ e = u

for u ∈ X^{∗},

x_{0}^{n} ⋆ w = w ⋆ x_{0}^{n} = wx_{0}^{n}

for any n ≥ 1 and w ∈ X^{∗}, and

(y_{s}u) ⋆ (y_{t}u′) = y_{s}(u ⋆ (y_{t}u′)) + y_{t}((y_{s}u) ⋆ u′) + y_{s+t}(u ⋆ u′)

for u and u′ in X^{∗}, s ≥ 1, t ≥ 1.

This product is extended by distributivity with respect to the addition to ℌ and defines a commutative and associative law. Moreover ℌ^{0} and ℌ^{1} are stable under ⋆. We denote by ℌ_{⋆}^{0} ⊂ ℌ_{⋆}^{1} ⊂ ℌ_{⋆} the corresponding harmonic algebras. Their structure has been investigated by M. Hoffman [H]: again they are (commutative) polynomials algebras over Lyndon words.

Specializing (4.1) at z_{1} = ⋯ = z_{k} = z′_{1} = ⋯ = z′_{k′} = 1, we deduce

(4.4)  ζ̂(u)ζ̂(u′) = ζ̂(u ⋆ u′)

for u and u′ in ℌ_{⋆}^{0}.

These are the second standard relations between multiple zeta values. For instance (4.3) with z = z′_{1} = z′_{2} = 1 gives

ζ(s)ζ(s′_{1}, s′_{2}) = ζ(s, s′_{1}, s′_{2}) + ζ(s′_{1}, s, s′_{2}) + ζ(s′_{1}, s′_{2}, s)

+ζ(s + s′_{1}, s′_{2}) + ζ(s′_{1}, s + s′_{2})

for s ≥ 2, s′_{1} ≥ 2 and s′_{2} ≥ 1.

5. The Third Standard Relations and the Main Diophantine Conjectures

We start with an example. Combining the stuffle relation (4.2) for s = s′ = 1 with the shuffle relation (2.2) for z′ = z, we deduce

(5.1) Li_{(1,2)}(z, 1) + 2Li_{(2,1)}(z, 1) = Li_{(1,2)}(z, z) + Li_{(2,1)}(z, z) + Li_{3}(z^{2}).

The two sides are analytic inside the unit circle, but not convergent at z = 1. We claim that

F (z) = Li_{(1,2)}(z, 1) − Li_{(1,2)}(z, z) = ∑_{n_{1}>n_{2}≥1} \frac{z^{n_{1}}(1 − z^{n_{2}})}{n_{1}n_{2}^{2}}

tends to 0 as z tends to 1 inside the unit circle. Indeed for |z| < 1 we have

|1 − z^{n_{2}}| = |(1 − z)(1 + z + ⋯ + z^{n_{2}−1})| < n_{2}|1 − z|,

hence

∑_{n_{2}=1}^{n_{1}−1} \frac{|1 − z^{n_{2}}|}{n_{2}^{2}} < |1 − z| ∑_{n_{2}=1}^{n_{1}−1} \frac{1}{n_{2}} ·

From (1.5) with n = 2 we deduce

|F (z)| ≤ |1 − z|Li_{(1,1)}(|z|) = \frac{1}{2}|1 − z|(log(1/(1 − |z|)))^{2}.

Therefore, taking the limit of the relation (5.1) as z → 1 yields Euler’s formula

ζ(2, 1) = ζ(3).

This argument works in a quite general setting and yields the relations

(5.2)  ζ̂(x_{1} ⋆ v − x_{1}� v) = 0

for each v ∈ ℌ^{0}.

These are the third standard relations between multiple zeta values. The Main Diophantine Conjectures below arose after the works of sev-

eral mathematicians, including D. Zagier, A.B. Goncharov, M. Kontsevich, M. Hoffman, M. Petitot and Hoang Ngoc Minh, K. Ihara and M. Kaneko (see [C]). They imply that the three standard relations (2.3), (4.4) and (5.2) generate the ideal of algebraic relations between all numbers ζ(\underline{s}). Here are precise statements.

We introduce independent variables Z_{u}, where u ranges over the set {e} ∪ X^{∗}x_{1}. For v = ∑_{u} c_{u}u in ℌ^{1}, we set

Z_{v} = ∑_{u} c_{u}Z_{u}.

In particular for u_{1} and u_{2} in x_{0}X^{∗}x_{1}, Z_{u_{1}� u_{2}} and Z_{u_{1}⋆u_{2}} are linear forms in Z_{u}, u ∈ x_{0}X^{∗}x_{1}. Also, for v ∈ x_{0}X^{∗}x_{1}, Z_{x_{1}� v−x_{1}⋆v} is a linear form in Z_{u}, u ∈ x_{0}X^{∗}x_{1}.

Denote by R the ring of polynomials with coefficients in ℚ in the variables Z_{u} where u ranges over the set x_{0}X^{∗}x_{1}, and by ℑ the ideal of R consisting of all polynomials which vanish under the specialization map

Z_{u} ↦ ζ̂(u) (u ∈ x_{0}X^{∗}x_{1}).

Conjecture 5.3. The polynomials

Z_{u_{1}}Z_{u_{2}} − Z_{u_{1}� u_{2}}, Z_{u_{1}}Z_{u_{2}} − Z_{u_{1}⋆u_{2}} and Z_{x_{1}⋆v−x_{1}� v},

where u_{1}, u_{2} and v range over the set of elements in x_{0}X^{∗}x_{1}, generate the ideal ℑ.

This statement is slightly different from the conjecture in § 3 of [IK], where Ihara and Kaneko suggest that all linear relations between MZV’s are supplied by the regularized double shuffle relations

ζ̂(reg(u ⋆ v − u� v)) = 0,

where u ranges over ℌ^{1} and v over ℌ^{0}. Here, reg is the ℚ-linear map ℌ → ℌ^{0} which maps w to the constant term of the expression of w as a (commutative) polynomial in x_{0} and x_{1} with coefficients in ℌ�^{0} . It is proved in [IK] that the linear polynomials Z_{reg(u⋆v−u� v)} associated to the regularized double shuffle relations belong to the ideal ℑ. On the other hand, at least for the small weights, one can check that the regularized double shuffle relations follow from the three standard relations. Hence our Conjecture 5.3 seems stronger than the Conjecture of [IK], but we expect they are in fact equivalent.

Denote by ℨ_{p} the ℚ-vector subspace of ℝ spanned by the real numbers ζ(\underline{s}) with \underline{s} of weight p, with ℨ_{0} = ℚ and ℨ_{1} = \{0\}. Using any of the first two standard relations (2.3) or (4.4), one deduces ℨ_{p} · ℨ_{p′} ⊂ ℨ_{p+p′}. This means that the ℚ-vector subspace ℨ of ℝ spanned by all ℨ_{p}, p ≥ 0, is a subalgebra of ℝ over ℚ which is graded by the weight. From Conjecture 5.3 one deduces the following conjecture of Goncharov [G1]:

Conjecture 5.4. As a ℚ-algebra, ℨ is the direct sum of ℨ_{p} for p ≥ 0.

Conjecture 5.4 reduces the problem of determining all algebraic relations between MZV to the problem of determining all linear such relations. The dimension d_{p} of ℨ_{p} satisfies d_{0} = 1, d_{1} = 0, d_{2} = d_{3} = 1. The expected value for d_{p} is given by a conjecture of Zagier [Z]:

Conjecture 5.5. For p ≥ 3 we have

d_{p} = d_{p−2} + d_{p−3}.

An interesting question is whether Conjecture 5.3 implies Conjecture 5.5. For this question as well as other related problems, see [É] and [C].

References

[BR] Ball, K., Rivoal, T. – Irrationalité d’une infinité de valeurs de la fonction zêta aux entiers impairs. Invent. Math., to appear.

[B^{3}] Borwein, J.M., Bradley, D.M., Broadhurst, D.J. – Evaluations of k-fold Euler/Zagier sums: a compendium of results for arbitrary k. Elec. J. Comb. 4 (1997), N^{∘} 2, # R5, 21 pp.

[B^{3}L] Borwein, J.M., Bradley, D.M., Broadhurst, D.J., Linosek, P. – Special Values of Multiple Polylogarithms. Trans. Amer. Math. Soc., 353 N^{∘} 3 (2001), 907-941.

[C] Cartier, P. – Fonctions polylogarithmes, nombres polyzêta et groupes prounipotents. Sém. Bourbaki, 53^{ème} année, 2000–2001, n^{∘} 884, Mars 2001, 36 pp.

[É] Écalle, J. – Rapport sur les multizêtas et les sommes d’Euler. Publ. Math. Orsay, 11 pp., to appear.

[Eu] Euler, L. – Meditationes circa singulare serierum genus. Leonhardi Euleri Opera Omnia, Series Prima XV, Commentationes Analyticae Vol. 2, 217–267; Novi Comm. Acad. Sci. Petropol., 20 (1775), 140-186.

[G1] Goncharov, A.B. – The double logarithms and Manin’s complex for modular curves. Math. Research Letter 4 (1997), n^{∘} 5, 6197–636.

[G2] Goncharov, A.B. – Multiple polylogarithms, cyclotomy and modular complexes. Math. Research Letter 5 (1998), 497–516.

[H] Hoffman, M.E. – The Algebra of Multiple Harmonic Series. J. Algebra 194 (1997) No.2, 477-495.

[IK] Ihara, K.; Kaneko, M. – Derivation relations and regularized double shuffle relations of multiple zeta values. Manuscript, 2001.

[K] Kassel, C. – Quantum Groups. Graduate Texts in Math. 155, Springer-Verlag,

1995.

[MPH] Minh H. N., Petitot, M., Van Der Hoeven J. – Shuffle algebra and polylogarithms. Proc. of FPSAC’98, 10-th international Conference on Formal Power Series and Algebraic Combinatorics, June 98, Toronto.

[R] Reutenauer, C. – Free Lie Algebras. London Math. Soc. Monographs New Series 7 (1993), Clarendon Press, Oxford.

[Ri1] Rivoal, T. – La fonction Zêta de Riemann prend une infinité de valeurs irrationnelles aux entiers impairs. C. R. Acad. Sci. Paris 331 (2000), 267–270.

http://arXiv.org/abs/math.NT/0008051

[Ri2] Rivoal, T. – Irrationalité d’au moins un des neuf nombres ζ(5), ζ(7), …, ζ(21). Submitted to Acta Arithmetica.

http://arXiv.org/abs/math/0104221

[W] Waldschmidt, M. – Valeurs zêta multiples: une introduction. J. Th. Nombres Bordeaux, 12 (2000), 581–595.

http://www.math.jussieu.fr/∼miw/articles/ps/MZV.ps

[Z] Zagier, D. – Values of zeta functions and their applications. First European Congress of Mathematics (Paris, 1992), Vol. II, Progr. Math. 120, Birkhäuser (1994) 497–512.

[Zu1] Zudilin, W. – Irrationality of values of zeta-function. To appear in the Proceedings of the Conference of Young Scientists (Moscow University, April 9-14, 2001)

http://arXiv.org/abs/math/0104249

[Zu2] Zudilin, W. – Arithmetics of linear forms involving odd zeta values. Manuscript, june 2001.

 http://www.math.jussieu.fr/∼miw/articles/ps/Chandigarh.ps